# Graviton-photon conversion in blazar jets as a probe of high-frequency gravitational waves

### v1

@@ I Introduction
A gravitational wave from a binary black hole merger was observed by LIGO in 2015 Abbott et al. (2016). This was the first observation of gravitational waves, which opened the era of the gravitational wave astronomy/cosmology. The frequency of the gravitational waves detectable with LIGO is around kHz. Pulsar timing arrays reported the evidence of the signal from stochastic gravitational waves in the frequency range of around nHz Agazie et al. (2023); Antoniadis et al. (2023). Therefore, multi-frequency gravitational wave observations have already begun, enabling us to explore new physics through the detection of various gravitational wave sources. In this regard, developing the methods for observing high-frequency gravitational waves above $10$kHz is demanded to probe interesting high-frequency gravitational wave sources Aggarwal et al. (2025); Ghiglieri and Laine (2015); Wang et al. (2019); Ghiglieri et al. (2020); Ringwald et al. (2021); Ghiglieri et al. (2022); Vagnozzi and Loeb (2022); Ema et al. (2022, 2015, 2016, 2020); Nakayama and Tang (2019); Huang and Yin (2019); Barman et al. (2023); Khlebnikov and Tkachev (1997); Easther and Lim (2006); Easther et al. (2007); Garcia-Bellido and Figueroa (2007); Ito and Soda (2016); Aggarwal et al. (2021); Franciolini et al. (2022); Saito et al. (2021); Gehrman et al. (2023a, b); Casalderrey-Solana et al. (2022); Ito and O. Ramos (2025).
@@
[…]
@@ II Mixing between gravitons and photons
$e^{+}_{i}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\enspace e^{\times}_{i}=\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},\enspace\epsilon^{+}_{ij}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix},\enspace\epsilon^{\times}_{ij}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}.$ (4)
Here, two linear polarizations of photons are leveled by $+$ and $\times$ according to plus and cross modes of gravitational waves. The electromagnetic field and the gravitational wave can be expanded as follows
$A_{i}$ $=e^{-i(wt-kz)}A^{\sigma}(z)e^{\sigma}_{i}\,,$ (5)
$h_{ij}$ $=e^{-i(wt-kz)}h^{\sigma}(z)\epsilon^{\sigma}_{ij}\,,$ (6)
@@
In deriving Eqs. (7) and (8), we assumed that the scale of graviton-photon conversion is much larger than $k^{-1}$, and that photons are ultrarelativistic, i.e., $\omega\simeq k$. We also neglected the spatial derivatives of the magnetic field $\bar{B}$. $T$ denotes the temperature of the cosmic microwave background. The plasma frequency $\omega_{p}$ and the cyclotron frequency $w_{c}$ are defined by
@@
where $\alpha$ is the fine structure constant, $n_{e}$ ($n_{p}$) is the electron (proton) number density, and $m_{e}$ ($m_{p}$) is the electron (proton) mass. $\gamma_{e}$ ($\gamma_{p}$) represents the Lorentz factor of electrons (protons), which suppresses the plasma frequencies since the effective masses of the plasma particles get heavier due to the Lorentz factors. Although the mixing between electromagnetic waves and gravitational waves is caused by the off-diagonal terms in Eqs. (7) and (8), the conversion efficiency is also affected by the diagonal terms, which correspond to the effective masses of the photons. Therefore, not only the magnitude and distribution of background magnetic fields, but also plasma properties such as number densities should be carefully investigated. As long as the magnitude of the magnetic field is not so large, we can ignore the contribution of the cyclotron frequency in Eq.(7) as $\omega_{c}/\omega<<1$ holds. Then, the plus and cross modes experience the same plasma effect approximately. Therefore, there is no actual difference in the conversion rates of the plus and cross modes in the low-frequency range where the plasma effect dominates the high order QED effect. On the other hand, in the high-frequency range where the high order QED effect is dominant, the conversion rates of plus and cross modes can become slightly different. In the next section, we study graviton-photon conversion in the magnetic fields of a jet by modeling the distributions of magnetic fields and plasma particles, with a particular focus on Mrk 501.
@@ III.1 Modeling of magnetic fields and plasma
Mrk 501 is a blazar located about 119 Mpc away ($z=0.034$), hosting a central black hole with a mass of $\sim 10^{9}$ solar mass of Technology . Although observations have not revealed the magnetic field configuration in the jet of Mrk 501 completely yet, it has been indicated that the helical component, which is almost orthogonal to jet direction, is dominant Hu et al. (2024) as is common for BL Lacertae objects Marscher and Jorstad (2021). This implies that graviton-photon conversion can occur effectively during the propagation within the blazar jet. In practice, an effective model of the magnetic field configuration is often employed to fit the spectral energy distribution (SED) of the target blazar. The one-zone synchrotron self-Compton (SSC) model is often adopted for this effective modeling Tavecchio et al. (1998). In the one-zone SSC model, relativistic charged particles emit synchrotron radiation and upscatter these photons through inverse Compton scattering, all within a single emission region of radius $r_{{\rm em}}$, which is filled with a tangled and homogeneous magnetic field of strength $B_{{\rm em}}$. Outside the emission region, one would expect that magnetic fields expand freely without radiative pressure. Thus, the strength of the magnetic field, envisioned as being dominated by its helical component, would scale as
$B(r)$ $=B_{{\rm em}}\hskip 64.01869pt\textrm{for}\enspace r\leq r_{{\rm em}},$ (11)
$B(r)$ $=B_{{\rm em}}\left(\frac{r}{r_{{\rm em}}}\right)^{-1}\quad\textrm{for}\enspace r_{{\rm em}}<r.$ (12)
Note that all quantities characterizing the jet models discussed here and in the following are measured in the comoving frame of the jet. We assume that the above scaling of the magnetic field continues to $r_{{\rm end}}$ at which its amplitude becomes comparable to the typical galactic magnetic fields, $\sim 1\mu$G. Below this field strength, the influence of the surrounding environment outside the jet may become significant, and the simple scaling may no longer be applicable.
@@
$N_{e}(\gamma)$ $=N_{0,e}\cdot\gamma^{-\alpha_{1,e}},\hskip 56.9055pt\textrm{for}\enspace\gamma_{{\rm min},e}\leq\gamma\leq\gamma_{{\rm br}}\,,$ (13)
$N_{e}(\gamma)$ $=N_{0,e}\cdot\gamma^{-\alpha_{2,e}}\gamma_{{\rm br}}^{\alpha_{2,e}-\alpha_{1,e}},\quad\textrm{for}\enspace\gamma_{{\rm br}}<\gamma\leq\gamma_{{\rm max},e}\,,$ (14)
@@
$N_{p}(\gamma)=N_{0,p}\cdot\gamma^{-\alpha_{p}},\quad\textrm{for}\enspace\gamma_{{\rm min},p}\leq\gamma\leq\gamma_{{\rm max},p}\,,$ (15)
@@
[…]
@@
The modeling of the jet configuration presented so far is general and not specific to Mrk 501. The parameter values in each model are determined to fit the SED of the blazar under study. In the next subsection, we focus on Mrk 501 and compute the graviton-photon conversion probability using the jet configuration consistent with its SED.
@@ III.2 Constraining gravitational waves with Mrk 501 observations
Requiring that produced photon flux within the bazar jet of Mrk 501 is smaller than the observed flux, one can give conservative constraints on the abundance of stochastic gravitational waves. The parameters characterizing Mrk 501 such as magnetic field strength, the number density of plasmas in the leptonic, lepto-hadronic, and hadronic one-zone SSC models are summarized in Table 1 based on Ref. Abe et al. (2023). The values of the magnetic fields already include the angular average of their components perpendicular to the gravitational wave propagation direction. We mention that the parameters are derived for the low-state SED of Mrk 501 observed during 2017-06-17 to 2019-07-23 (MJD 57921 to MJD 58687), so that the contribution of transitional high states Acciari et al. (2011); Aliu et al. (2016); Ahnen et al. (2018); Sahu et al. (2020); Acciari et al. (2020) has been excluded Abe et al. (2023).
@@
$r_{{\rm em}}$[cm] $1.14\times 10^{17}$ $1.14\times 10^{17}$ $1.14\times 10^{17}$
@@
Figure 1: Upper limits on $h_{c}$ of stochastic gravitational waves are shown. The top-left, top-right, and bottom panels correspond to the hadronic, leptonic, and lepto-hadronic models, respectively. The green solid and dashed lines represent conservative and aggressive constraints, respectively, derived by considering intergalactic magnetic fields Ito et al. (2024b). Conservative and aggressive constraints from ARCADE2 and EDGES, based on CMB spectral distortions due to graviton-photon conversion, are also shown Domcke and Garcia-Cely (2021). Limits from axion search experiments are indicated as CAST, OSQAR, and ALPS Ejlli et al. (2019).
Figure 2: Upper limits on $h^{2}\Omega_{{\rm GW}}$ of stochastic gravitational waves are shown. The top-left, top-right, and bottom panels correspond to the hadronic, leptonic, and lepto-hadronic models, respectively. The green solid and dashed lines represent conservative and aggressive constraints, respectively, derived by considering intergalactic magnetic fields Ito et al. (2024b). Conservative and aggressive constraints from ARCADE2 and EDGES, based on CMB spectral distortions due to graviton-photon conversion, are also shown Domcke and Garcia-Cely (2021). Limits from axion search experiments are indicated as CAST, OSQAR, and ALPS Ejlli et al. (2019).

### v2

@@ I Introduction
A gravitational wave from a binary black hole merger was observed by LIGO in 2015 Abbott et al. (2016). This was the first observation of gravitational waves, which opened the era of the gravitational wave astronomy/cosmology. The frequency of the gravitational waves detectable with LIGO is around kHz. Pulsar timing arrays reported the evidence of the signal from stochastic gravitational waves in the frequency range of around nHz Agazie et al. (2023); Antoniadis et al. (2023). Therefore, multi-frequency gravitational wave observations have already begun, enabling us to explore new physics through the detection of various gravitational wave sources. In this regard, developing the methods for observing high-frequency gravitational waves above 1010kHz is demanded to probe interesting high-frequency gravitational wave sources Aggarwal et al. (2025); Ghiglieri and Laine (2015); Wang et al. (2019); Ghiglieri et al. (2020); Ringwald et al. (2021); Ghiglieri et al. (2022); Vagnozzi and Loeb (2022); Ema et al. (2022, 2015, 2016, 2020); Nakayama and Tang (2019); Huang and Yin (2019); Barman et al. (2023); Khlebnikov and Tkachev (1997); Easther and Lim (2006); Easther et al. (2007); Garcia-Bellido and Figueroa (2007); Ito and Soda (2016); Aggarwal et al. (2021); Franciolini et al. (2022); Saito et al. (2021); Gehrman et al. (2023a, b); Casalderrey-Solana et al. (2022); Ito and O. Ramos (2025); Dong et al. (2016).
@@
Another way to probe high-frequency gravitational waves is to utilize the electromagnetic wave probe, whose typical frequency range is $10^{8}$$\sim 10^{35}Hz$, with the use of graviton-photon conversion Pshirkov and Baskaran (2009); Dolgov and Ejlli (2012); Domcke and Garcia-Cely (2021); Ramazanov et al. (2023); Liu et al. (2024); Ito et al. (2024a, b); Dandoy et al. (2024); He et al. (2024); Lella et al. (2024); McDonald and Ellis (2024); Kushwaha and Jain (2025); Li et al. (2025). Graviton-photon conversion inevitably occurs in the presence of a background magnetic field. It implies that gravitational waves in the universe convert to electromagnetic fields universally, since various magnetic fields exist across a wide range of scales such as intergalactic magnetic fields, galactic magnetic fields, and magnetic fields around stars. Possibilities for probing high-frequency gravitational waves with graviton-photon conversion in the intergalactic magnetic fields Domcke and Garcia-Cely (2021); Ito et al. (2024b); He et al. (2024); Kushwaha and Jain (2025); Li et al. (2025), galactic magnetic fields Ito et al. (2024b); Lella et al. (2024), pulsar magnetic fieldsIto et al. (2024a); Dandoy et al. (2024); McDonald and Ellis (2024), and geomagnetic fields Liu et al. (2024); Ito et al. (2024b) have been studied so far. In this paper, we study graviton-photon conversion within the magnetic fields of a blazar and give new constraints on stochastic gravitational waves by using observational data of Mrk 501.
@@ II Mixing between gravitons and photons
$e^{+}_{i}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\kern 5.0pte^{\times}_{i}=\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},\kern 5.0pt\epsilon^{+}_{ij}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix},\kern 5.0pt\epsilon^{\times}_{ij}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$. (4)
Here, two linear polarizations of photons are leveled by ++ and $\times$ according to plus and cross modes of gravitational waves. The electromagnetic field and the gravitational wave can be expanded as follows
$A_{i}$ $=e^{-i(\omega t-kz)}A^{\sigma}(z)e^{\sigma}_{i}\,$, (5)
$h_{ij}$ $=e^{-i(\omega t-kz)}h^{\sigma}(z)\epsilon^{\sigma}_{ij}\,$, (6)
@@
In deriving Eqs. (7) and (8), we assumed that the scale of graviton-photon conversion is much larger than $k^{-1}$, and that photons are ultrarelativistic, i.e., $\omega\simeq k$. We also neglected the spatial derivatives of the magnetic field $\bar{B}$. $T$ denotes the temperature of the cosmic microwave background. The plasma frequency $\omega_{p}$ and the cyclotron frequency $\omega_{c}$ are defined by
@@
where $\alpha$ is the fine structure constant, $n_{e}$ ($n_{p}$) is the electron (proton) number density, and $m_{e}$ ($m_{p}$) is the electron (proton) mass. $\gamma_{e}$ ($\gamma_{p}$) represents the Lorentz factor of electrons (protons), which suppresses the plasma frequencies since the effective masses of the plasma particles get heavier due to the Lorentz factors. Although the mixing between electromagnetic waves and gravitational waves is caused by the off-diagonal terms in Eqs. (7) and (8), the conversion efficiency is also affected by the diagonal terms, which correspond to the effective masses of the photons. Therefore, not only the magnitude and distribution of background magnetic fields, but also plasma properties such as number densities should be carefully investigated. As long as the magnitude of the magnetic field is not so large, we can ignore the contribution of the cyclotron frequency in Eq.(7) as $\omega_{c}/\omega\ll 1$ holds. Then, the plus and cross modes experience the same plasma effect approximately. Therefore, there is no actual difference in the conversion rates of the plus and cross modes in the low-frequency range where the plasma effect dominates the high order QED effect. On the other hand, in the high-frequency range where the high order QED effect is dominant, the conversion rates of plus and cross modes can become slightly different. In the next section, we study graviton-photon conversion in the magnetic fields of a jet by modeling the distributions of magnetic fields and plasma particles, with a particular focus on Mrk 501.
@@ III.1 Modeling of magnetic fields and plasma
Mrk 501 is a blazar located about 119 Mpc away ($z=0.034$), hosting a central black hole with a mass of $\sim 10^{9}$ solar mass of Technology . Although observations have not revealed the magnetic field configuration in the jet of Mrk 501 completely yet, it has been indicated that the helical component, which is almost orthogonal to jet direction, is dominant Hu et al. (2024) as is common for BL Lacertae objects Marscher and Jorstad (2021). This implies that graviton-photon conversion can occur effectively during the propagation within the blazar jet. In practice, an effective model of the magnetic field configuration is often employed to fit the spectral energy distribution (SED) of the target blazar. The one-zone synchrotron self-Compton (SSC) model is often adopted for this effective modeling Tavecchio et al. (1998). In the one-zone SSC model, relativistic charged particles emit synchrotron radiation and upscatter these photons through inverse Compton scattering, all within a single emission region of radius $r_{{\rm em}}$, which is filled with a tangled and homogeneous magnetic field of strength $B_{{\rm em}}$, meaning that the magnitude and directional dependence are averaged, in the jet’s comoving frame. Outside the emission region, one would expect that magnetic fields expand freely without radiative pressure. Thus, the strength of the magnetic field, envisioned as being dominated by its helical component, would scale as
$B(r)$ $=B_{{\rm em}}\hskip 64.01869pt\textrm{for}\kern 5.0ptr\leq r_{{\rm em}},$ (11)
$B(r)$ $=B_{{\rm em}}\left(\frac{r}{r_{{\rm em}}}\right)^{-1}\quad\textrm{for}\kern 5.0ptr_{{\rm em}}<r.$ (12)
[…]
@@
$N_{e}(\gamma)$ $=N_{0,e}\cdot\gamma^{-\alpha_{1,e}},\hskip 56.9055pt\textrm{for}\kern 5.0pt\gamma_{{\rm min},e}\leq\gamma\leq\gamma_{{\rm br}}\,$, (13)
$N_{e}(\gamma)$ $=N_{0,e}\cdot\gamma^{-\alpha_{2,e}}\gamma_{{\rm br}}^{\alpha_{2,e}-\alpha_{1,e}},\quad\textrm{for}\kern 5.0pt\gamma_{{\rm br}}<\gamma\leq\gamma_{{\rm max},e}\,$, (14)
@@
$N_{p}(\gamma)=N_{0,p}\cdot\gamma^{-\alpha_{p}},\quad\textrm{for}\kern 5.0pt\gamma_{{\rm min},p}\leq\gamma\leq\gamma_{{\rm max},p}\,$, (15)
@@
$\bar{\gamma}_{i}=\frac{\int\gamma N_{i}(\gamma)d\gamma}{\int N_{i}(\gamma)d\gamma}.$ (16)
@@
The modeling of the jet configuration presented so far is general and not specific to Mrk 501. The parameter values in each model are determined to fit the SED of the blazar under study.
In the next subsection, we focus on Mrk 501 and compute the graviton–photon conversion probability using the one-zone SSC modeling of the jet configuration consistent with its SED. This treatment allows us to reduce the problem to one dimension, as a consequence of averaging the magnetic field distribution in the jet’s comoving frame. We believe that this approximation is valid for estimating the order of magnitude. Moreover, a more effective graviton–photon conversion would be expected if the helical component is dominant in an actual jet, which may indeed be favored Hu et al. (2024); Marscher and Jorstad (2021). Therefore, compared to the case without angular averaging of the magnetic field distribution, the present one-zone modeling may provide a conservative estimate of the graviton–photon conversion probability for leptonic, lepto-hadronic, and hadronic models.
@@ III.2 Constraining gravitational waves with Mrk 501 observations
Requiring that produced photon flux within the blazar jet of Mrk 501 is smaller than the observed flux, one can give conservative constraints on the abundance of stochastic gravitational waves. The parameters characterizing Mrk 501 such as magnetic field strength, the number density of plasmas in the leptonic, lepto-hadronic, and hadronic one-zone SSC models are summarized in Table 1 based on Ref. Abe et al. (2023). The values of the magnetic fields already include the angular average of their components perpendicular to the gravitational wave propagation direction. We mention that the parameters are derived for the low-state SED of Mrk 501 observed during 2017-06-17 to 2019-07-23 (MJD 57921 to MJD 58687), so that the contribution of transitional high states Acciari et al. (2011); Aliu et al. (2016); Ahnen et al. (2018); Sahu et al. (2020); Acciari et al. (2020) has been excluded Abe et al. (2023).
@@
$r_{{\rm em}}$[cm] $14\times 10^{17} 1.1$$4\times 10^{17} 1.1$$14\times 10^{17}$
@@
Figure 1: Upper limits on $h_{c}$ of stochastic gravitational waves are shown. The top-left, top-right, and bottom-left panels correspond to the hadronic, leptonic, and lepto-hadronic models, respectively. The bottom-right panel shows a direct comparison among the three models. The green solid and dashed lines represent conservative and aggressive constraints, respectively, derived by considering intergalactic magnetic fields Ito et al. (2024b). Conservative and aggressive constraints from ARCADE2 and EDGES, based on CMB spectral distortions due to graviton-photon conversion, are also shown Domcke and Garcia-Cely (2021). Limits from axion search experiments are indicated as CAST, OSQAR, and ALPS Ejlli et al. (2019).
Figure 2: Upper limits on $h^{2}\Omega_{{\rm GW}}$ of stochastic gravitational waves are shown. The top-left, top-right, and bottom-left panels correspond to the hadronic, leptonic, and lepto-hadronic models, respectively. The bottom-right panel shows a direct comparison among the three models. The green solid and dashed lines represent conservative and aggressive constraints, respectively, derived by considering intergalactic magnetic fields Ito et al. (2024b). Conservative and aggressive constraints from ARCADE2 and EDGES, based on CMB spectral distortions due to graviton-photon conversion, are also shown Domcke and Garcia-Cely (2021). Limits from axion search experiments are indicated as CAST, OSQAR, and ALPS Ejlli et al. (2019).